### v2

@@ 2 Airdrop Desiderata
dYdX Sept. 2021 Still active Ethereum Continuous DEX
1inch Dec. 2020 Still active Ethereum Continuous DEX
[…]
Arbitrum Mar. 2023 Sept. 2023 Arbitrum One Single Single
Uniswap Sept. 2020 Still active Ethereum Continuous DEX
[…]
Next, we break down airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals. These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
Foster long-term involvement.  Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points. For example, in layer-2 blockchains (L2s) [67], one may award discounts on future transactions. Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the blockchain has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of their various features and use-cases [44].
Target value-creating users.  An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [78] and DEXs [77], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
3 Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them [38], and 61% have sold their tokens shortly after the drop in the case of ParaSwap [21].
In both instances, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely sybil accounts. Additionally, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. Here, we examine data associated with six airdrops (ENS, dYdX, 1inch, Arbitrum, and Uniswap, including Gemstone) presented in Table 1, which were gathered from archive nodes of Ethereum and zkSync Era.To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query: select * from addresses_ethereum.dex and from Etherscan [20].
Table 2: General airdrop distribution statistics.
[…]
supply distributed exchanges addresses
[…]
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$
1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$
Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$
[…]
Table 3: Tokens received per address.
[…]
dYdX $2325.85$ $91336.75$ $392.56$ 00 $24.7\times 10^{6}$
1inch $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Arbitrum $1874.02$ $1566.92$ $1250$ $625$ $10250$
Uniswap $620.13$ $9347.91$ $400$ $400$ $2.1\times 10^{6}$
[…]
3.1 Tokens Distributed Over Time
Table 2 presents the number of airdrops distributed and subsequently sold on exchanges. Additionally, Table 3 details the distribution of airdrops per user. These tables provide quantitative insights into how airdropped tokens are disseminated across the user base. Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2). Furthermore, the data indicate that tokens were frequently traded on exchanges. This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2). Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges. In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [35].
Figure 1: Tokens claimed per day for each of the Airdrops analyzed.
Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users. Figure 1 illustrates the distribution of accounts claiming their tokens each day. Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts. Notably, $583137$ (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens. Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day. This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.
Table 4: Airdrop recipients’ interaction with exchanges.
[…]
Total # sent to exchanges % sent to exchanges
[…]
dYdX $42682$ $41320$ 96.81%
1inch $43402$ $41913$ 96.57%
Gemstone $20014$ $20000$ 99.93%
Arbitrum $583137$ $555040$ 95.18%
[…]
Table 5: Distribution of hops for airdrop recipients connecting with exchanges.
[…]
Mean Std. Median Min. Max.
[…]
dYdX 2.25 0.59 2 1 18
1inch 1.85 0.64 2 1 7
Gemstone 1 1 1 1 1
Arbitrum 2.71 $0.79$ $3$ $1$ $32$
Uniswap 2.11 $0.43$ 2 1 9
[…]
(a) Over the number of post-airdrop blocks.
(b) Over the number of post-airdrop days.
Figure 2: Time for airdrop tokens to be transferred to exchanges.
Additionally, Table 5 presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset. Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities. Gemstone is a notable outlier, with all tokens being sold within a single hop. Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (between 2 and 3 hops). This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
@@ 3.1 Tokens Distributed Over Time
Figure 2 shows the cumulative distribution functions (CDFs) demonstrating how quickly users interact with an exchange right after receiving tokens from an airdrop. As shown in Figure 2(a), the majority of accounts exchanged their tokens within approximately one million blocks. The number of blocks for Gemstone is significantly higher than others due to a delay introduced by airdrop farmers. Given the difference in block generation times between Ethereum (with a new block created approximately every 15 seconds) and zkSync, we normalized the blocks in terms of days. According to Figure 2(b), within a single day, 64.67% of the accounts interacted with an exchange. In comparison, ENS had an interaction rate of 45.86%, while dYdX stood at 37.41%. This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2. The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
3.2 Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges. Lastly, Arbitrum accounts for $921650$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Figure 3. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [20]. We show that for all protocols but Gemstone, the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.

### v3

@@ 4.3 Total Value Locked (TVL)
5 Common Airdrop Design Challenges
Airdrops, much like traditional loyalty programs—such as newcomer bonuses provided by banks and credit card companies—face several common design challenges. However, the unique context of blockchain technology and the specific mechanics employed by most airdrops can amplify these challenges or even introduce new ones. In this section, we explore three of these challenges.
5.1 Airdrop Farmers
These are sophisticated users who employ elaborate tactics to maximize the amounts of airdrop tokens they receive. Blockchain protocols have adopted various measures to mitigate airdrop farmers gaming the system. A popular approach is to limit the amount of rewards that a single user can receive.
Thus, protocols resort to Proof-of-Humanity (PoH)services, such as Gitcoin passport [28]. These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human. Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [28]. Such methods can be augmented by analyzing on-chain data to detect Sybils and exclude them from the airdrop [5], but this may lead to false negatives [69]. There are some recent examples of protocols like Celestia using Sybil resistance solutions [40].
There are other mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [56], to sharing posts on social media [38]. These tasks sometimes appear arbitrary, leading to frustration among users and they are easy to automate [34] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [55]. Furthermore, the reliance of many protocols on a limited number of PoHservices means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full Sybil resistance is not guaranteed [65]. A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement. Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [9], as was the case in dYdX’s airdrop [13].
5.2 Decentralized Governance Threats
Some protocols distribute governance tokens through airdrops as a means to decentralize their governance processes. However, distributing governance tokens can be risky. These tokens allow holders to participate in the protocol’s governance by voting on key decisions. Typically, these tokens can also be exchanged for other tokens, potentially giving them monetary value [12]. This can lead to more farmers farming these tokens.
Empirical evidence suggests that airdropping governance tokens may be more effective than those of non-governance tokens. In a recent analysis, Makridis et al. [45] found that airdropped governance tokens experienced a growth rate in market capitalization up to $14.99$% higher compared to governance tokens that were not airdropped. However, the authors also note that this effect is statistically insignificant when using common benchmarks.
Despite these potential benefits, airdropping governance tokens poses significant risks if not done well. It can concentrate too much power in the hands of a few users, leading to an inequitable distribution of decision-making power within the system [17, 48, 60]. Additionally, some recipients may not have the protocol’s best interests in mind and could act in ways detrimental to the protocol’s long-term success by voting on changes to the protocol that benefit themselves.
5.3 Insider Trading
This issue arises when individuals exploit privileged information for financial gain at the expense of other protocol users. Such practices are widely recognized as violations of securities laws in traditional financial markets [4] and often provoke negative reactions from the blockchain community [18, 52, 61].
When someone within a protocol uses privileged information to increase their profits, it can trigger a community backlash. Insiders may have advanced knowledge of the metrics used to determine eligibility and the reward for each address, and may use it to their advantage. For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200$K USD profit from their airdrop, although later dismissed as a mere coincidence [61]. However, such incidents can undermine the trust users place in these protocols.
This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others. Identifying these inside traders is a challenging problem, making it crucial for protocols to provide extensive information to their users. In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [18]. For this to happen, data availability is key. Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
6 Design Guidelines
The design challenges outlined earlier, while sobering, may inform future airdrop designers, and can shed light on the path to potential success.
6.1 Alternative Incentives for Sustaining User Engagement
The long-term benefits of an airdrop to a protocol can be difficult to measure, as the potential advantages may be indirect and hard to quantify, while the expenses are often immediate and irrevocable. Additionally, some costs and effects, such as those arising from distributing governance tokens, may be unpredictable. Therefore, instead of using airdrops, communities could consider alternative measures with a more predictable relationship between costs and benefits. One straightforward alternative is for the community to vote to programmatically reward loyal users with discounts on future interactions. In the context of Layer 2 (L2) solutions, these discounts could be applied to transaction fees. This approach encourages users to engage with the protocol again to benefit from the incentive, thereby fostering continued user participation. Additionally, this incentive mechanism is somewhat more resistant to airdrop farming, as the discounts have no intrinsic value outside the protocol, and the protocol’s costs are directed solely toward users who actively use the system.
However, careful design of the discount mechanism is essential, including determining eligibility criteria for users and setting appropriate discount levels. Moreover, it is not immediately clear whether discounts can attract users as effectively as the more immediate and tangible rewards offered by airdrops. Another option is to perform multiple airdrops over an extended period instead of a single event. Although this approach remains susceptible to some pitfalls of standard one-shot airdrops, it may help ensure long-term community engagement and prevent the decline in user adoption that some protocols experience immediately after an airdrop. A more innovative approach was adopted by Blast, which introduced a point-based reward program [47]. Under this program, users earn rewards based on points accumulated through various activities, such as bridging tokens to the protocol (i.e., transferring funds from another protocol) and participating in a referral program that increases points for bringing in more users. Notably, Blast received \$1.1 billion in deposits even before its official launch [36]. This method provides a measurable metric for users’ contributions to the protocol, structured around a referral-program model.
Furthermore, the incorporation of innovative distribution mechanisms plays a crucial role in mitigating the presence of sybil attacks within whitelisted addresses eligible for an airdrop. For example, Celestia has proposed a unique design that uses GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [22]. However, a potential concern arises with the possibility of users or farmers generating counterfeit activities on GitHub to exploit other protocols’ airdrops adopting a similar strategy to Celestia’s. Farmers, therefore, might expect that new protocols use similar selection criteria than past airdrops. To counter this, protocols could focus on metrics that are resistant to programmable manipulation, thereby adding a layer of difficulty or cost to the creation of automated user accounts.
6.2 Targeting Well-Known Reputable Entities
Instead of rewarding pseudonymous users, protocols can target developers and projects that are building relevant applications. For example, in Arbitrum’s airdrop, 1.13% of the distributed tokens were allocated to DAOs projects [19]. Arbitrum also offers additional incentives beyond its airdrop for specific groups, such as university students [21] and members of the technical community [20] who wish to participate in research and develop tools relevant to the protocol.
Another approach was implemented by Optimism, which allocated some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup to the blockchain world [23]. Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can promote sustained engagement. By funding these entities, protocols may attract value-driven users and foster long-term involvement.
6.3 Proactive Oversight and Community Engagement
During the airdrop process, continuous monitoring and analysis of the protocol’s data are crucial to prevent exploitation. A notable example is provided by Linea, whose team identified a vulnerability that allowed users to manipulate their incentive mechanism [41]. This prompt discovery averted cheaters from claiming over one-third of the Non-Fungible Tokens (NFTs)offered as incentives [58].
In addition to technical oversight, engaging the community by maintaining open communication channels and offering bug bounties encourages the disclosure of vulnerabilities, whether exploited or not [64]. For instance, an AzukiDAO community member disclosed a vulnerability, allowing the protocol to quickly address it [49]. Oversight should also extend beyond on-chain data. For example, social media is frequently exploited by scammers promoting fake airdrops, luring users into connecting their wallets to fraudulent websites that drain their funds. Even protocols not planning an airdrop may be targeted by such schemes.
Active community involvement in technical discussions is also advantageous. For example, ZKsync Era’s NFTdrop was retrospectively analyzed by cygaar [11], who identified potential cost-saving improvements. Maintaining transparency and providing the community with insights into the protocol’s inner workings fosters trust. When technical issues arise, users are more likely to respond with understanding when they are well-informed.
6.4 Rewards Should Scale with Costs
The effects of Goodhart’s law [46] are evident in many past airdrops. For example, airdrops often aim to incentivize user engagement by publicly announcing that eligibility for rewards is based on the frequency of past interactions with the protocol. This approach, however, can be exploited, as users may engage in meaningless wash-trading transactions [67], undermining the criteria as a true reflection of genuine engagement.
The problem is further exacerbated by the fact that action-based metrics used to determine eligibility often fail to consider the actual costs users incur for each action. For instance, when the number of transactions is the primary metric, low transaction fees enable airdrop farmers to cheaply meet the transaction volume required for rewards. A potential solution is the adoption of a reputation-based mechanism [32], which could disincentivize artificial inflation of transaction volumes. However, the protocol must carefully define what constitutes “user reputation” and determine the appropriate metrics for evaluation.
Conversely, high transaction fees can erode the value of the rewards users receive, reducing the attractiveness of the airdrop [51]. To address these issues, rewards should scale with the actual costs incurred by users, ensuring a fairer and more effective distribution of incentives.
7 Related Work
Recent research on airdrops has primarily focused on post-factum analyses and guidance for designing effective airdrop campaigns.
Airdrop Studies  Yaish and Livshits [71] present a theoretical model for airdrops that accounts for both honest users and farmers, where the latter have lower airdrop eligibility costs and lower intrinsic utility from using the platform that issues the airdrop. In their analysis, they show that airdrops in which the issuer expends a non-zero fixed cost for each recipient, the threat posed by farmer sybil attacks lead to unbounded issuance costs. However, they show losses from farmers can be bounded by setting in advance the total quantity of airdrop tokens and splitting them equally among recipients. Moreover, by designing the airdrop mechanism correctly, farmers can be harnessed to foster network effects, thereby attracting honest users who may otherwise would prefer competing platforms.
Makridis et al. [45] explore the impact of governance token airdrops on DEXgrowth. Analyzing 51 exchanges, they found that such airdrops significantly contribute to increased market capitalization and trading volume. Lommers et al. [44] provide a comprehensive overview of various airdrop types, such as plain vanilla, holder-based, and value-based models. Their work highlights how eligibility criteria, signaling, and implementation strategies influence the success of airdrop campaigns, offering practical recommendations for optimization. Fan et al. [16] present a case study of ParaSwap DEX, proposing a role taxonomy methodology based on user behavior and airdrop effectiveness. Their findings show that users receiving higher rewards are more likely to contribute positively to the community. They also identify arbitrage patterns and note the limitations of current methods for detecting airdrop hunters. On the other hand, Sybil detection techniques, such as graph network analysis [16, 43] and machine learning approaches [8, 31, 50, 66, 73], have been proposed to address these issues.
Allen [2] conducts nine airdrop case studies (e.g., Optimism, Arbitrum, Blur) and offers insights into designs like task-based claiming. The study emphasizes the need for more dynamic designs with feedback loops and suggests that some projects may revert to simpler designs due to the complexity and cost of advanced mechanisms. Allen et al. [3] investigate the motivations behind token airdrops, focusing on marketing and decentralization. While airdrops are often seen as marketing tools, the authors argue that this rationale is weak, as evidence of marketing-driven airdrop success is limited. Instead, decentralization and community-building are highlighted as the primary motivations.
Technical Aspects of Airdrops  Frowis et al. [26] identify the operational challenges and costs associated with large-scale airdrops on Ethereum. They suggest that while cost savings of up to 50% are possible with specific smart contract optimizations, pull-based approaches can shift the cost burden to recipients. Overall, the total cost remains proportional to the number of recipients.
Wahby et al. [68] address the privacy issues in current airdrop mechanisms, which leak information about recipients. They propose a private airdrop scheme based on zero-knowledge proofs using RSA credentials, achieving privacy while maintaining computational efficiency. Their implementation significantly improves both signature generation and verification times.
8 Conclusion
This study identified common issues with airdrops and proposed guidelines to improve their effectiveness. Our analysis of six major airdrops revealed a consistent pattern of recipients quickly selling their tokens—36.62% of ENS, 35.45% of dYdX, and 54.05% of 1inch tokens were traded shortly after distribution, with a median of just two transactions. This suggests airdrops fail to sustain long-term engagement or attract valuable contributors.
For Arbitrum, we observed a spike in daily fees during the airdrop, followed by a decline in transactions per address. Other protocols outperformed Arbitrum without an airdrop, and since June 2023, transaction fees have been similar across protocols, indicating airdrops are not the main driver of user growth.
Lastly, we discussed challenges like airdrop farming, governance token distribution, and insider trading, offering insights for future airdrop strategies.
Acknowledgments
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
Appendix 0.A Reproducibility
To ensure full reproducibility of our empirical analysis, we will openly share both our dataset and scripts. The primary data source used in §3 consists of blockchain data extracted from Ethereum, Arbitrum, and ZKsync Era archive nodes. We collected all relevant blocks, transactions, and events tied to the contracts of the six airdrops examined.
For §4, data was sourced from Growthepie.xyz [29] corresponding to the year 2023, which offers various metrics, including daily transaction counts, daily active addresses, median transaction costs, TVL, user fees, and stablecoin market capitalization. These metrics are available for major blockchains such as ZKsync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM. To enhance clarity, all metrics were smoothed using a 7-day moving average.
Appendix 0.B Graph Network of the Token Distribution
We construct a graph network (V, E) of the token transfer network for the six airdrops analyzed in this paper and present them in Fig. 7. Each node (V) represents an address, and an edge (E) is created whenever tokens are transferred from one address to another.
For the ENS network, there are $184585$ nodes and $608462$ edges. The dYdX network consists of $112853$ nodes and $406027$ edges. The Gemstone network contains $20003$ nodes and $240113$ edges. The 1inch network includes $308329$ nodes and $1400913$ edges. The Arbitrum network has $921650$ nodes and $27438608$ edges, while the Uniswap network contains $1180830$ nodes and $3762613$ edges.
@@ Appendix 0.B Graph Network of the Token Distribution
Appendix 0.C More Insights Into Airdrop Metrics
In this section, we provide a complementary analysis of the measures described in Section §4. To summarize, Fig. 8 shows the ratio of unique daily active addresses between Arbitrum and Optimism. Additionally, Fig. 9 presents the total daily transaction fees before and after the Arbitrum airdrop.
(a) Absolute daily transaction fees.
(b) Relative to average.
Figure 9: Total daily transaction fees before and after the Arbitrum airdrop: (a) absolute values of daily fees in USD; and (b) fees relative to the average before the airdrop.
Appendix 0.D Top Airdrop Earners
In this section, we present the top-1 airdrop earners for each of the six protocols analyzed. Airdrops can be distributed to various types of accounts, including smart contract accounts, such as the protocol’s own treasury, or Externally Owned Accounts (EOAs). Additionally, tokens can be sent to multi-signature (multisig) contract wallets. A multisig contract requires a minimum number of users to sign a transaction before it can be executed on-chain [48]. These multisig accounts may also belong to the protocol itself. For instance, the treasury account could receive a substantial portion of the airdropped tokens, which could be used to support the protocol’s ecosystem. Therefore, the accounts who receive most of the tokens (the top earners) may be protocol-related accounts (see Table 2 in §3.1).
To identify the top airdrop earner for each protocol, we list them in Table 4. The labels for these earners were obtained using the methodology detailed in §3.2, which involves gathering data from Etherscan, a widely-used blockchain explorer [15]. We note that analytico.eth alone received $11.71$% of the 1inch airdropped tokens. The true identity of the entity controlling this address remains unknown, and it is unclear whether it is managed by a single individual or a group. On the other hand, $13.95$% of dYdX were sent to dYdX treasure. This is an expected behavior for some airdrops.
Table 4: Top airdrop earner for the six protocols analyzes in this paper: ENS, dYdX, 1inch, Gemstone, Arbitrum, and Uniswap.
Protocol Wallet address Label # of tokens % of total supply % of dist. tokens
ENS 0xa5b5$\cdots$6c2d bentobox.eth $1143.54$ $0.0011$ $0.006$
dYdX 0x95$\cdots$0b54 TreasuryMerkleClaimProxy $24.7\times 10^{6}$ $2.47$ $13.95$
1inch 0xa0f7$\cdots$9b73 analytico.eth $9.75\times 10^{6}$ $0.65$ 11.71
Gemstone 0$\cdots$3a3f Sybil [30] $2.64\times 10^{12}$ $5.21\times 10^{-39}$ $0.005$